OpenStudy (anonymous):

How many ounces of pure cranberry juice and how many ounces of a citrus fruit drink containing 10% fruit juice should be mixed to get 630 oz of a fruit drink that is 23% fruit juice?

OpenStudy (anonymous):

x=pure y=citrus 630=x+y 630*.23=x+.1y 144.9=x+.1(630-x) 144.9=x+63-.1x 0.9x=81.9 x=91 AWNS